OpenStudy (anonymous):

use the quadratic formula to solve the equation 3x^2-7x-5=0

OpenStudy (anonymous):

Ready to work on this?

OpenStudy (anonymous):

Lemme know if you needed help soon. Unless you just wanted an answer, in which case, nvm.

OpenStudy (anonymous):

3x^2-7x-5=0 calculate the discriminant D = b² - 4ac D = (-7)² - 4(3)(-5) = 49 + 60 = 109 - b +/- root(b² - 4ac) -(-7) +/- root(109) x = --------------------- = ------------------- 2a 2(3) 7 + root(109) 7 - root(109) so x = -------------- or -------------- 6 6

OpenStudy (anonymous):

@Harkirat: The problem is he/she is most likely going to only see the part where you wrote x = blah blah and put that into the online homework and be content. It's more important that we verify that they actually understand it.

OpenStudy (anonymous):

yeah, u r right to an extent... only a foolish student will just copy, the ones who want to learn come back with questions and get things ceared and then try on their own n ask me to check...

OpenStudy (anonymous):

Indeed. But I shouldn't really care about it I guess. No reason in me trying to help someone who just wants answers. So nvm about what I wrote. Sorry. :/

OpenStudy (anonymous):

hey, u said nothing tht u need to apologize for...... good eggs, bad eggs.....a shopkeeper has to handle all !!!! ☺